Tags:
create new tag
view all tags
<!-- keep this as a security measure:
* Set ALLOWTOPICCHANGE = TWikiAdminGroup,Main.LCGAdminGroup,Main.EgiGroup
* Set ALLOWTOPICRENAME = TWikiAdminGroup,Main.LCGAdminGroup
#uncomment this if you want the page only be viewable by the internal people
#* Set ALLOWTOPICVIEW = TWikiAdminGroup,Main.LCGAdminGroup,Main.ChippComputingBoardGroup
-->

Swiss Grid Operations Meeting on 2018-04-05 at 14:00

Site status

CSCS

System

- 57 nodes on LHConCray

- Close to a solution for CMS on LHConCray: With a compiled version of condor for Suse and also some minor script changes to gliteWMS (https://github.com/holzman/glideinWMS/tree/master/creation/web_base)

- Working on ARC for Cern Project

Storage

GPFS
- New pool "cache" activated for small/most used files. Now metadata and most used files are on High End Flash System
- Huge performance improvement on the storage servers and Phoenix WN (see attachment "wn-performance-march2018.png")
- Some tuning to do on DVS nodes (high context switch issue)
- In the next weeks will receive new All-SSD storage system for CHiPP and CERN


dCache
- New servers will be delivered in the next weeks
- ATLAS data dump recovered

PSI

UNIBE-LHEP

    • Stable operation for several months, no issues or immediate worries to report.
    • Running an average of 2500 slots, Ubelix contribution 23%

    • Accounting numbers (from scheduler) from last month

      VOJob TypeProduced WC core-hours
      ATLAS Any 1341763 (was 1184331)
      ops Any 12 (was 20)
      t2k.org Any

      11104 (was 72)

      uboone Any 0 (was 0)


      • Accounting numbers (from dashboard) from last month for CSCS and UNIBE

      • HC availability [1]:
        • CSCS-LCG2: 95% Prod (was 97%), 95% Analy (was 97%)
        • CSCS-LCG2-HPC: 80% (was 91%) Prod, 82% (was 91%) Analy
        • UNIBE-LHEP: 100% (was 98%) Prod, 100% (was 95%) Analy
        • UNIBE-LHEP-UBELIX: 100% (was 98%) Prod, 100% (was 97%) Analy

          * CSCS running 2680 (was 3000) slots on average, UNIBE running 2500 (was 2400)


Cluster Job Type Produced WC core-hours Good vs Bad WC % CPU eff good jobs %
CSCS Any 2103505 (was 2373724) (53%) 0.87 (was 0.62) 0.83 (was 0.58)
Unibe Any 1736362 (was 1507416) (47%) 0.89 (was 0.94) 0.86 (was 0.80)






[1] http://dashb-atlas-ssb.cern.ch/dashboard/request.py/siteviewhistorywithstatistics?columnid=562#time=custom&start_date=2018-03-01&end_date=2018-03-31&use_downtimes=false&merge_colors=false&sites=multiple&clouds=all&site=ANALY_CSCS,ANALY_CSCS-HPC,ANALY_UNIBE-LHEP,ANALY_UNIBE-LHEP-UBELIX,CSCS-LCG2,CSCS-LCG2-HPC,CSCS-LCG2-HPC_MCORE,CSCS-LCG2_MCORE,UNIBE-LHEP,UNIBE-LHEP-UBELIX,UNIBE-LHEP-UBELIX_MCORE,UNIBE-LHEP_MCORE

UNIBE-ID

  • Stable delivery for ATLAS (~600 slots continuously)
  • Still planning to implement the event service to fetch oppurtunistically idle slots. No progress yet (lack ot time)

UNIGE

  • Site still in extended downtime
  • The DPM SE has been decommissioned
  • The disk servers are being re-installed and expected to be attached to the Bern DPM head node (first DPM distributed setup)
  • The ARC CE is expected to be re-deployed right after that

NGI_CH

* NGI-CH Open Tickets review

Ticket-ID Type VO Site Priority Resp. Unit Status Last Update Subject Scope
134196   cms CSCS-LCG2 urgent NGI_CH assigned 2018-03-27 Update Phedex download agent parameters WLCG
133885   none   less urgent NGI_CH in progress 2018-03-14 NGI_CH - February 2018 - RP/RC OLA ... EGI
133787 lhcb CSCS-LCG2 very urgent NGI_CH waiting for reply 2018-04-05 Pilots Failed at CSCS-LCG2 HPC WLCG
133695 lhcb CSCS-LCG2 urgent NGI_CH waiting for reply 2018-04-05 Data access problem at CSCS-LCG2 WLCG
132927   cms CSCS-LCG2 urgent NGI_CH involved in progress 2018-03-01 Problem with APEL Accounting for all of ... EGI
131965   none UNIBE-LHEP less urgent NGI_CH on hold 2018-03-19 IPv6 deployment at WLCG Tier-2 sites EGI
131948   none CSCS-LCG2 less urgent NGI_CH assigned assigned 2018-03-01 IPv6 deployment at WLCG Tier-2 sites EGI
131433   none T3_CH_PSI less urgent NGI_CH assigned in progress 2018-03-12 Storage accounting deployment EGI

Other topics

  • Topic1
  • Topic2

    Next meeting date:

A.O.B.

Attendants

  • CSCS:
  • CMS:
  • ATLAS:
  • LHCb:
  • EGI:

Action items

  • Item1
Edit | Attach | Watch | Print version | History: r4 < r3 < r2 < r1 | Backlinks | Raw View | Raw edit | More topic actions
Topic revision: r4 - 2018-04-05 - DinoConciatore
 
This site is powered by the TWiki collaboration platform Powered by Perl This site is powered by the TWiki collaboration platformCopyright © 2008-2018 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback